### v1

# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

Abstract
Inferring trajectories from longitudinal spatially-resolved omics data is fundamental to understanding the dynamics of structural and functional tissue changes in development, regeneration and repair, disease progression, and response to treatment. We propose ContextFlow, a novel context-aware flow matching framework that incorporates prior knowledge to guide the inference of structural tissue dynamics from spatially resolved omics data. Specifically, ContextFlow integrates local tissue organization and ligand-receptor communication patterns into a transition plausibility matrix that regularizes the optimal transport objective. By embedding these contextual constraints, ContextFlow generates trajectories that are not only statistically consistent but also biologically meaningful, making it a generalizable framework for modeling spatiotemporal dynamics from longitudinal, spatially resolved omics data. Evaluated on three datasets, ContextFlow consistently outperforms state-of-the-art flow matching methods across multiple quantitative and qualitative metrics of inference accuracy and biological coherence. Our code is available at: https://github.com/santanurathod/ContextFlow
1 Introduction
Flow matching (Lipman et al., 2023) is an emerging paradigm that provides an efficient approach for learning the complex latent dynamics, or normalizing flows (Papamakarios et al., 2021), of a system of variables, while enabling parametric flexibility to model data distributions. Inferring the underlying dynamics from sparse and noisy observations is a central challenge in many domains (Gontis et al., 2010; Brunton et al., 2016; Pandarinath et al., 2018; Li et al., 2025), where continuous trajectories are rarely captured; instead, cross-sectional snapshots, collected at discrete time points, are typically available. In single-cell RNA sequencing (scRNA-seq), this challenge becomes especially critical as the destructive nature of profiling technologies yields only unpaired population-level snapshots over time. Uncovering temporal dynamics from such snapshot data is essential for understanding developmental processes, disease progression, treatment and perturbation responses (Wagner & Klein, 2020). Traditional approaches often rely on heuristics or computationally intensive likelihood-based generative models, which struggle with scalability and flexibility in high-dimensional single-cell data. Flow matching overcomes these challenges by directly learning continuous latent dynamics that are constrained to match observed population-level distributions at sampled time points.
The state and function of cells within a tissue are affected by interactions with neighboring cells, extracellular matrix components, and local signaling gradients (Rao et al., 2021). Recent advances in spatial omics technologies, particularly spatial transcriptomics (ST), allow gene expression profiling without tissue dissociation, thereby preserving spatial context and providing a complementary view of cellular organization. The dynamics of complex cellular processes is affected by the tissue microenvironment, where cells engage in reciprocal communication with their neighbors (Dimitrov et al., 2022; Tanevski et al., 2025). A growing body of work highlights the critical role of spatial cell–cell communication patterns in shaping cellular phenotypes (Armingol et al., 2021). In particular, location-specific communication circuits between distinct cell types dynamically interact to reprogram cellular states and influence tissue-level behavior (Mayer et al., 2023; Aguadé-Gorgorió et al., 2024; Zheng et al., 2025). These insights, made possible by the spatiotemporal resolution of transcriptomics data, pave the way for understanding the mechanisms by which cellular interactions drive tissue organization and function in organogenesis (Chen et al., 2022), regeneration (Ben-Moshe et al., 2021; Wei et al., 2022), disease progression (Kukanja et al., 2024), and treatment response (Liu et al., 2024).
Optimal transport (OT) has become a foundational framework to align spatially resolved samples and infer putative developmental or temporal couplings  (Zeira et al., 2022; Liu et al., 2023). As a result, state-of-the-art flow matching frameworks such as minibatch-OT flow matching (MOTFM) (Tong et al., 2024) use OT-derived couplings to define conditional paths to train velocity fields, thus overcoming the lack of generative capabilities in optimal transport. The OT formulation adopted in MOTFM, however, does not account for the contextual richness present in spatial transcriptomics and can result in trajectories that are statistically optimal yet biologically implausible (see Figure 3(a) in Appendix G.1 for an illustration). While recent studies have extended widely-used OT objectives (Halmos et al., 2025; Ceccarelli et al., 2025) for spatial transcriptomics, they primarily focus on pairwise alignment of populations across conditions or modalities and do not explicitly incorporate the cell–cell communication patterns that drive cellular state transitions.
To address the above limitations, we introduce a novel flow matching-based framework, ContextFlow, that incorporates spatial priors for modeling temporal tissue dynamics (Figure 1). By encoding local tissue organization and ligand-receptor-derived spatial communication patterns into prior-regularized optimal transport formulations, ContextFlow fully exploits the contextual richness of spatial omics data and embeds both structural and functional aspects of tissue organization into its objective, thereby generating more biologically informed trajectories. In summary, our contributions are as follows:
  • We leverage local tissue organization and local ligand–receptor communication patterns to extract biologically meaningful features from spatial omics data, and encode them into a biologically-informed transition plausibility matrix to constrain temporal dynamics (Section 3.2).
  • We design two novel integration schemes—cost-based and entropy-based—that incorporate the prior knowledge into an OT-coupled flow matching framework, both amenable to efficient Sinkhorn optimization and scalable on modern hardware (Section 3.3).
  • Comprehensive experiments on regeneration and developmental datasets demonstrate that ContextFlow consistently outperforms baseline methods under both interpolation and extrapolation settings across metrics that capture biological plausibility and statistical fidelity (Section 4).
Figure 1: ContextFlow integrates local tissue organization and ligand-receptor communications to learn biologically meaningful trajectories from spatial omics data. Prior knowledge acts as a soft filter that discourages implausible transitions while preserving flexibility in trajectory inference.
2 Preliminaries
2.1 Flow Matching Basics
Flow matching (Lipman et al., 2023) is a simulation-free and sample-efficient generative framework for training continuous normalizing flows (Chen et al., 2018). Given a pair of source and target data distributions over $\mathbb{R}^{d}$ with densities $q_{0}=q(\bm{x}_{0})$ and $q_{1}=q(\bm{x}_{1})$, the problem task is to learn a time-varying velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, whose continuous evolution is captured by a function in the form of a neural-net-based model with weights $\theta$, that can transform $q_{0}$ to $q_{1}$ through integration via an ordinary differential equation (ODE). To be more specific, flow matching (FM) seeks to optimize $\theta$ by minimizing a simple regression loss between $u_{\theta}$ and a target time-varying velocity vector field $u_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ as follows:
$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{x}\sim p_{t}(\bm{x})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x})\big\|^{2}.$ (1)
Here, $\mathcal{U}(0,1)$ is the uniform distribution over $[0,1]$, and $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ denotes a time-varying probability path induced by $u_{t}$ such that (i) $p_{t}$ is a probability density function for any $t\in[0,1]$, (ii) $p_{t}$ satisfies the two boundary conditions: $p_{t=0}=q_{0}$ and $p_{t=1}=q_{1}$, and (iii) the connection between $p_{t}$ and $u_{t}$ can be characterized by the transport equation (Villani et al., 2008): $\frac{\partial p_{t}(\bm{x})}{\partial t}=-\nabla\cdot(u_{t}(\bm{x})p_{t}(\bm{x}))$, where $\nabla$ is the divergence operator. From a dynamical system’s view, $u_{t}$ defines an ODE system $d\bm{x}=u_{t}(\bm{x})dt$. The corresponding solution to the ODE, usually termed as the probability flow, can then transport any $\bm{x}_{0}\sim q_{0}$ to a point $\bm{x}_{1}\sim q_{1}$ along $u_{t}$ from $t=0$ to $t=1$. While the flow matching objective in Equation 1 is simple and intuitive, it is generally intractable in practice: the closed-form velocity vector field $u_{t}$ is unknown for arbitrary source and target distributions ($q_{0}$ and $q_{1}$), and multiple valid probability paths $p_{t}$ may exist between them.
2.2 Conditional Flow Matching
The central idea of conditional flow matching is to express the target probability path $p_{t}$ via a mixture of more manageable conditional probability paths (Lipman et al., 2023). By marginalizing over some conditioning variable $z$, both $p_{t}$ and $u_{t}$ can be constructed using their conditional counterparts:
$p_{t}(\bm{x})=\int p_{t}(\bm{x}|\bm{z})q(\bm{z})d\bm{z},\quad u_{t}(\bm{x})=\int u_{t}(\bm{x}|\bm{z})\frac{p_{t}(\bm{x}|\bm{z})q(\bm{z})}{p_{t}(\bm{x})}d\bm{z},$ (2)
where $q(\bm{z})$ denotes the distribution of the conditioning variable $\bm{z}$, and $p_{t}(\bm{x}|\bm{z})$ is selected such that the boundary conditions are satisfied: $\int p_{t=0}(\bm{x}|\bm{z})q(\bm{z})=q_{0}$ and $\int p_{t=1}(\bm{x}|\bm{z})q(\bm{z})=q_{1}$. Theorem 1 of Lipman et al. (2023) proves that $p_{t}$ and $u_{t}$ defined by Equation 2 satisfy the transport equation, suggesting that $p_{t}$ is a valid probability path generated by $u_{t}$. To avoid the intractable integrals, Lipman et al. (2023) proposed the following optimization of conditional flow matching (CFM), and proved its equivalence to the original flow matching objective in terms of gradient computation:
$\min_{\theta}\>\mathbb{E}_{t\sim\mathcal{U}(0,1),z\sim q(\bm{z}),\bm{x}\sim p_{t}(\bm{x}|\bm{z})}\big\|u_{\theta}(t,\bm{x})-u_{t}(\bm{x}|\bm{z})\big\|^{2}.$ (3)
By choosing an appropriate conditional velocity vector field $u_{t}(\bm{x}|\bm{z})$, we can train the neural network using Equation 3 without requiring a closed-form solution of the conditional probability path $p_{t}(\bm{x}|\bm{z})$, thus avoiding the intractable integration operation. Therefore, the remaining task is to define the conditional probability path and velocity vector field properly such that we can sample from $p_{t}(\bm{x}|\bm{z})$ and compute $u_{t}(\bm{x}|\bm{z})$ efficiently for solving the optimization problem in Equation 3.
Gaussian Conditional Probability Paths. A specific choice proposed in Lipman et al. (2023) is Gaussian conditional probability paths and their corresponding conditional velocity vector fields:
$p_{t}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x}\>|\>\mu_{t}(\bm{z}),\sigma_{t}(\bm{z})^{2}\mathbf{I}),\quad u_{t}(\bm{x}|\bm{z})=\frac{\sigma_{t}^{\prime}(\bm{z})}{\sigma_{t}(\bm{z})}(\bm{z}-\mu_{t}(\bm{z}))+\mu_{t}^{\prime}(\bm{z}),$ (4)
where $\mu_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the time-varying mean of the Gaussian distribution, $\mu_{t}^{\prime}$ is its derivative with respect to time, $\sigma_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{+}$ stands for the time-varying scalar standard deviation, and $\sigma_{t}^{\prime}$ stand for the corresponding derivative. In particular, Lipman et al. (2023) set $q(\bm{z})=q(\bm{x}_{1})$, $\mu_{t}(\bm{z})=t\bm{x}_{1}$, and $\sigma_{t}(\bm{z})=1-(1-\sigma)t$. Then, we can see that $u_{t}(\bm{x}|\bm{z})$ transports the standard Gaussian distribution $p_{t=0}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};0,I)$ to a Gaussian distribution with mean $x_{1}$ and standard deviation $\sigma$, namely $p_{t=1}(\bm{x}|\bm{z})=\mathcal{N}(\bm{x};\bm{x}_{1},\sigma^{2})$ for any target point $\bm{x}_{1}$. By letting $\sigma\rightarrow 0$, the marginal boundary conditions can easily be verified. Tong et al. (2024) further generalized the application scope to arbitrary source distributions, by setting
$q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1}),\quad\mu_{t}(\bm{z})=(1-t)\bm{x}_{0}+t\bm{x}_{1},\quad\sigma_{t}(\bm{z})=\sigma.$ (5)
This choice satisfies the boundary conditions $p_{t=0}(\bm{x})=q_{0}$ and $p_{t=1}(\bm{x})=q_{1}$ when $\sigma\rightarrow 0$. Based on Equation 4, the conditional velocity vector field has a simple analytical form $u_{t}(\bm{x}|\bm{z})=\bm{x}_{1}-\bm{x}_{0}$.
2.3 Flow Matching with Optimal Transport Couplings
The conditionals construction specified by Equation 5 corresponds to the simplest choice of independent coupling, where $\bm{z}=(\bm{x}_{0},\bm{x}_{1})$ with source $\bm{x}_{0}$ and target $\bm{x}_{1}$ are independently sampled from $q(\bm{z})=q(\bm{x}_{0})q(\bm{x}_{1})$. The use of couplings for constructing the sampling paths in the CFM framework naturally connects to the optimal transport theory (Villani et al., 2008). Choosing OT-based couplings has several advantages over independent coupling, including smaller training variance and more efficient sampling (Pooladian et al., 2023; Tong et al., 2024).
Since the classical Kantorovich’s formulation (refer Appendix E) has computational complexity that is cubic with respect to the sample size, a popular alternative is to add an extra regularization term, resulting in entropic optimal transport (EOT), to approximately solve the optimal transport problem while reducing the computational costs from cubic to quadratic:
$\pi^{*}_{\mathrm{eot}}(\epsilon):={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1})+\epsilon H(\pi\>|\>q_{0}\otimes q_{1}),$ (6)
where $\epsilon>0$ is the regularization parameter, and $H(\pi\>|\>q_{0}\otimes q_{1})$ denotes the relative entropy (or Kullback-Leibler divergence) with respect to $\pi$ and the product measure $q_{0}\otimes q_{1}$. The optimization problem in Equation 6 can be viewed as a special case of the static Schrödinger bridge problem (Bernton et al., 2022), which can be efficiently solved in a mini-batch fashion via the Sinkhorn algorithm (Cuturi, 2013). Theoretically, one can prove that $\pi^{*}_{\mathrm{eot}}(\epsilon)$ recovers the Kantorovich’s OT coupling $\pi^{*}_{\mathrm{ot}}$ when $\epsilon\rightarrow 0$ (see Equation 16 in Appendix E for its formal definition) and $\pi^{*}_{\mathrm{eot}}(\epsilon)$ corresponds to the independent coupling $q_{0}\otimes q_{1}$ when $\epsilon\rightarrow\infty$.
3 Regularizing the Flow with Spatial Priors
3.1 Problem Formulation
We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data. Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points. For simplicity, we use $[m]$ to denote the set $\{1,2,\ldots,m\}$. For any $i\in[m+1]$, let $q_{i}$ be the data distribution at time $t_{i}$ over $\mathbb{R}^{d}$. Given $\{\mathbf{X}_{t_{i}}\}_{i\in[m+1]}$, where $\mathbf{X}_{t_{i}}=[\bm{x}_{i}(k)]_{k\in[n_{i}]}$ is the expression matrix at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$. This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points. Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, define
$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),\quad u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (7)
where the conditioning variable is selected as $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 6. It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$. To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and integrate the biological prior knowledge that can be inferred from other associated data modalities. Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension. However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 3(a) in Appendix G.1 for an illustration).
3.2 Introducing Spatial Priors & Transitional Plausibility
To faithfully model the spatial context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the transitional plausibility between locations and cell states at different time points.
Spatial Smoothness. Tissues are well-organized systems. Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affects their states in a similar manner and results in local smoothness of cell-type-specific expression. Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale. However, the same heterogeneity allows us to consider the spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2025) as a proxy for relative cell localization, which cannot change significantly across short time intervals. Therefore the aggregate expression within the microenvironment of each cell can be used to quantify the transitional plausibility in consecutive time points.
Specifically, let $c_{i}=(\bm{x}_{i},\bm{s}_{i})$ and $c_{j}=(\bm{x}_{j},\bm{s}_{j})$ be cells at time points $t_{i}$ and $t_{j}$, respectively, where $\bm{x}_{i},\bm{x}_{j}\in\mathbb{R}^{d}$ denote their gene expression profiles, and $\bm{s}_{i},\bm{s}_{j}\in\mathbb{R}^{2}$ denote their spatial coordinates in the relative tissue reference frame. Let $\mathrm{TP}(c_{i},c_{j})$ denote the transitional plausibility, i.e., the likelihood that $c_{i}$ evolves to $c_{j}$ between $t_{i}$ and $t_{j}$. Spatial smoothness suggests that $\mathrm{TP}(c_{i},c_{j})$ is inversely related to the difference between the average expression profiles of their local neighborhoods:
$\mathrm{SS}(c_{i},c_{j})=\bigg\|\>\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\>\bigg\|_{2}^{2},$ (8)
where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ is the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile of cell $c$.
Cell-Cell Communication Patterns. Cell–cell communication (CCC) has a critical role in the regulation of numerous biological processes, including development, apoptosis, and the maintenance of homeostasis in health and disease (Armingol et al., 2024). A major type of CCC is ligand–receptor (LR) signaling, in which ligands expressed by one cell bind to cognate receptors on another, initiating intracellular cascades that ultimately affect the state of the cell (i.e., its expression profile)  (Armingol et al., 2021). There are numerous databases of prior knowledge of ligand-receptor binding and computational methods that use these databases to systematically link gene expression with the activity of ligand-receptor-mediated communication.
Specifically, we can represent each cell $c_{i}$ by a vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair. The $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 7 for an illustration). We define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
$\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2},$ (9)
3.3 ContextFlow: CFM with Context-Aware OT Couplings
Our proposed framework, graphically depicted in Figure 1, consists of the following three main steps:
Transitional Plausibility Matrix. First, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points. Specifically, let $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ be the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}\left(c_{i}(k),c_{i+1}(l)\right)+(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right),$ (10)
where $\lambda\in[0,1]$ is a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
Prior-Regularized OT Couplings. The transitional plausibility matrices capture our spatially informed prior on cell-cell transitions between consecutive time points, which can naturally be incorporated in the EOT formulation (Equation 6) to promote couplings that maintain the structural and functional properties of the tissue organization. We propose two techniques for prior integration:
Prior-Aware Cost Matrix (PACM). Consider the empirical counterpart of Equation 6 with respect to time $t_{i}$ and time $t_{i+1}$. Our first approach incorporates the transitional plausibility matrix directly into the transport cost:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\underbrace{\Big[\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot\big[\mathbf{M}_{i,i+1}\big]_{kl}\Big]}_{\text{Prior-Aware Cost Function}}-\epsilon\sum_{k,l}\Pi_{kl}\log\Pi_{kl},$ (11)
where the transport plan $\Pi$ satisfies the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility. If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 11 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$. This aligns with our assumption that such transitions are implausible.
Prior-Aware Entropy Regularization (PAER). While the prior-aware cost matrix approach penalizes couplings in accordance with our spatial priors, it defines a different OT problem characterized by a modified cost function. Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds. Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison. This normalization, however, may result in couplings that deviate from their original counterparts (Proposition 1 and Corollary 1 in the Appendix C). Besides, selecting an appropriate $\alpha$ in Equation 11 introduces an additional layer of tuning, increasing computational overhead. Therefore, we propose a second approach to integrate the biological priors without introducing additional hyperparameters:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}$ $\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\log(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl})}_{\text{Prior-Aware Entropy Regularization}},$ (12)
where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\exp(-[\mathbf{M}_{i,i+1}]_{kl})/\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})$ denotes the prior joint probability matrix induced by $\mathbf{M}_{i,i+1}$. Intuitively, the lower the cost $\mathbf{M}{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}{i,i+1}]_{kl}$, reflecting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$. The entropy regularization term in Equation 12 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
ContextFlow. Finally, we apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problem in Equation 11 or Equation 12 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 7. The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings (ContextFlow), is detailed in Algorithm 1 in Appendix D.
In particular, to apply the Sinkhorn algorithm to solve our prior-aware entropy regularization problem in Equation 12, we make use of the following theorem, a generalized result of Peyré et al. (2019).
Theorem 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix. Suppose $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
$\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}={\arg\!\min}_{\mathbf{\Pi}\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
where $\epsilon>0$ is the regularization parameter. Then, we can show that $\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}$ can be computed by Sinkhorn and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ denotes element-wise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
Theorem 1, proven in Appendix B, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices. When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 16. When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with Entropic-OT (Section 2.3). This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow. By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.
4 Experiments
Datasets. We evaluate ContextFlow on three longitudinal spatial transcriptomics datasets: Axolotl Brain Regeneration (Wei et al., 2022), Mouse Embryo Organogenesis (Chen et al., 2022), and Liver Regeneration (Ben-Moshe et al., 2021). For all the datasets, the gene expression values are log-normalized, and we extract the top $50$ principal components (PCs) as feature vectors. The strength of ligand–receptor interactions in the microenvironment was inferred using spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), where we applied the cosine similarity metric to gene expression profiles. Interaction evidence was aggregated using the consensus of multiple curated ligand–receptor resources, ensuring robustness of the inferred signals.
Baselines & Metrics. We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several baselines. As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings. We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors. For evaluation, we employed 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, and metrics such as MMD and Energy Distance for statistical fidelity. Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset. Exact metric definitions are present in the Appendix F. All reported metrics are averaged across $10$ runs.
Sampling. A trained velocity field can be evaluated through the samples it generates. We consider two variants. Initial value problem sampling (IVP) integrates the learned gradient starting from the first observed batch of cells and evolves them toward a later time point. IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps. In contrast, next-step sampling (Next Step) integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
Table 1: Interpolation at the middle holdout time point for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
CTF-C 11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
CTF-C 11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
4.1 Axolotl Brain Regeneration
We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022). The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage. For our CTF-C method, we present the best ablated $\alpha$ in the main text, with full ablation results across different $\alpha$ values provided in Appendix H.
For interpolation, we hold out the middle time point during training and evaluate it using samples generated by the trained velocity field $u_{\theta}$ via both IVP and next-step sampling. Table 1 presents the results. Across multiple evaluation metrics, ContextFlow with entropy regularization (CTF-H) produces trajectories that most closely match the ground truth. CTF-H consistently achieves the best or comparable performance relative to CTF-C, despite the latter being explicitly tuned across multiple $\alpha$ values. This highlights the computational efficiency and superior generalization ability of CTF-H, as it avoids the need for additional hyperparameter tuning while maintaining strong performance.
For extrapolation, we evaluate generation on the last holdout time point, representing the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon. As shown in Table 2, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate. This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon. Finally, Figure 3 (Appendix G.1) demonstrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.
Table 2: Extrapolation on the last holdout time point for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.275\pm 0.011$ $76.947\pm 5.661$
MOTFM $619\pm 0.611 7.7$$769\pm 0.763$ $0.272\pm 0.007$ $85.352\pm 8.140$
CTF-C 11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
CTF-C 11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
Table 3: Interpolation (time 5) and extrapolation (time 8) results on the Organogenesis dataset.
Method $\bm{\lambda}$ $\bm{\alpha}$ Next Step (Interpolation) IVP (Interpolation) Next Step (Extrapolation)
Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$
MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $3.251\pm 0.676$ $3.418\pm 0.727$ $1.626\pm 0.066$ $1.682\pm 0.096$
CTF-C 11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $3.137\pm 0.407$ $4.093\pm 1.187$ $1.685\pm 0.096$ $1.714\pm 0.160$
0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $2.938\pm 0.476$ $3.904\pm 1.120$ $1.773\pm 0.053$ $1.880\pm 0.180$
$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $3.200\pm 0.403$ $3.555\pm 0.637$ $1.768\pm 0.058$ $1.858\pm 0.120$
11 $0.2$ $1.880\pm 0.020$ $1.922\pm 0.078$ $3.260\pm 0.880$ $5.264\pm 3.060$ $1.683\pm 0.058$ $1.803\pm 0.117$
0 $0.2$ $1.900\pm 0.035$ $1.912\pm 0.057$ $2.953\pm 0.425$ $3.816\pm 0.970$ $1.715\pm 0.123$ $1.860\pm 0.267$
CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $3.244\pm 0.713$ $3.946\pm 1.671$ $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$
11 $1.898\pm 0.029$ $1.866\pm 0.097$ $5.200\pm 0.799$ $6.306\pm 1.037$ $1.890\pm 0.046$ $1.877\pm 0.103$
$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $1.636\pm 0.060$ $1.684\pm 0.099$
(a) Brain Regeneration
(b) Organogenesis
Figure 2: KL-Divergence between predicted and ground-truth cell type distributions.
Table 4: Interpolation results on the middle holdout time point for the Liver Regeneration dataset.
MOTFM CTF-C CTF-H
($\lambda$, $\alpha$) (11, $0.5$) (0, $0.5$) ($0.5$, $0.8$) (0, –) (11, –) ($0.5$, –)
$\bm{\mathcal{W}_{2}}$ $4.303\pm 1.448 33$$3.506\pm 1.148$ $2.741\pm 1.864 33$$3.045\pm 1.644$ $\bm{32.682\pm 1.472}$ $3.481\pm 1.001 33$$3.414\pm 0.995$
4.2 Mouse Embryo Organogenesis
We further evaluated ContextFlow on the larger Mouse Organogenesis Spatiotemporal Atlas (MOSTA) Stereo-seq dataset (Chen et al., 2022) spanning measurements from $8$ developmental time points. For the interpolation study of this dataset, we held out time point 5 during training and evaluated its generation during testing. Table 3 shows the evaluation results. We observe that ContextFlow, with both integration strategies, outperforms MOTFM across all metrics, showcasing the effectiveness of the contextual information. While CTF-C shows stronger performance under next-step sampling—albeit only after fine-tuning the trade-off parameter $\alpha$—CTF-H consistently outperforms it in the more challenging IVP-Sampling setting. On the extrapolation task, integrating to the final time point, CTF-H again achieves the strongest performance, underscoring that the entropy-regularized formulation not only removes the need for additional parameter tuning but also offers more robust generalization to unseen temporal horizons.
Figure 2 reports the KL-Divergence between normalized histograms of predicted and ground-truth cell types from ContextFlow and MOTFM. In both cases, CTF exhibits lower divergence on average across time points, indicating that the trajectories generated by our model better preserve the biological composition of cell types over time. The cell type progression is further visualized in Figure 9 (Appendix H.6). We show that ContextFlow predicts temporal cell type trajectories that evolve smoothly and consistently across consecutive developmental stages. Early progenitor populations, such as neural crest and mesenchyme, progressively diminish as development advances, while terminal fates, including muscle, cartilage primordium, and liver, emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout, demonstrating that ContextFlow captures biologically coherent and temporally consistent developmental dynamics.
4.3 Liver Regeneration
Finally, we evaluate ContextFlow on a Visium spatial transcriptomics dataset profiling the temporal dynamics of mouse liver regeneration following acetaminophen-induced injury (Ben-Moshe et al., 2021), collected across three distinct regeneration stages. Unlike the earlier datasets resolved at single-cell resolution, Visium data is captured at the level of $55$ micron diameter spots, capturing the joint expression of multiple cells. Since direct cell-type information is unavailable, we restrict evaluation to the 2-Wasserstein distance. Moreover, since evaluation is performed on the middle of the three time points, IVP and next-step predictions coincide. Table 4 presents the results. Consistent with the previous findings, CTF-H achieved the lowest reconstruction error, indicating that incorporating contextual information improves trajectory estimation even in aggregated spot-level measurements.
5 Conclusion
We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data, addressing a central challenge of existing methods. The entropic variant of ContextFlow is theoretically grounded, which always yields OT couplings constrained by prior knowledge, promoting stability and consistency with the imposed contextual constraints. Across three diverse datasets, we showed that ContextFlow consistently improves over state-of-the-art baselines even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors. In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and results in coherent and temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics. These results highlight the value of embedding biological context into generative flow models. Future works can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or leverage multi-marginal OT formulations for optimizing temporal flows. Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, bridging generative power with biological interpretability.
Appendix A Related Work
A.1 Flow Matching
Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021). These transformations can be realized through either finite (Rezende & Mohamed, 2015) or continuous compositions (Chen et al., 2018). The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive. Flow matching (FM) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable. To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024). Owing to this scalability, FM has been rapidly adopted across scientific domains, including biology and the life sciences (Li et al., 2025). In transcriptomics, for example, Klein et al. (2024) employed an FM backbone to approximate OT maps for drug response modeling and cross-modal translation tasks. Entropic OT formulations have also been applied to infer cellular trajectories (Tong et al., 2024; Rohbeck et al., 2025), generate imaging-based cell morphology changes (Zhang et al., 2025), and simulate spatial transcriptomics data from histology images (Huang et al., 2025).
Despite these advances, existing work does not address how to meaningfully incorporate biological prior knowledge to constrain the velocity field, limiting the biological plausibility of inferred trajectories.
A.2 Optimal Transport
Omics studies frequently generate uncoupled measurements across conditions, modalities, or time points, which must be integrated into a unified representation to provide a more comprehensive view of the underlying biology. Optimal transport (OT) has recently gained popularity for this task, as it provides a geometry-based approach to couple probability distributions (Bunne et al., 2024; Klein et al., 2025). In spatial transcriptomics (ST), several OT formulations have been introduced depending on context. For instance, Zeira et al. (2022) and Liu et al. (2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation. Rahimi et al. (2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al. (2025) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer.
While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
Appendix B Proofs of Main Theoretical Results
Proposition 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then, for indices $(i,j)$ and $(k,l)$,
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
where $\gamma$ depends on $\Pi^{*}_{ij},c$ and the OT marginal constraints $a,b$.
Proof.
For the original optimal transport (OT) formulation, we note:
$\Pi^{*}_{ij}=u_{i}K_{ij}v_{j},\quad K_{ij}=e^{-C_{ij}/\epsilon},$
with the constraints $\Pi^{*}\mathbf{1}=a$ and $\Pi^{*\top}\mathbf{1}=b$.
Let
$\Pi_{ij}^{*1/c}=u_{i}^{1/c}K_{ij}^{1/c}v_{j}^{1/c},$
where:
$\tilde{K}_{ij}=K_{ij}^{1/c}=\exp\left(-C_{ij}/(c\epsilon)\right)$
is the kernel for the scaled/normalized OT formulation. Let $\tilde{\Pi}^{*}_{ij}$ be the coupling for the scaled version, then:
$\tilde{\Pi}^{*}_{ij}=\tilde{u}_{i}\tilde{K}_{ij}\tilde{v}_{j}.$
Thus, there exist scaling factors $\alpha_{i},\beta_{j}\in\mathbb{R}$ such that:
$\tilde{u}_{i}$ $=\alpha_{i}u^{\frac{1}{c}}_{i},$
$\tilde{v}_{j}$ $=\beta_{j}v^{\frac{1}{c}}_{j}.$
This implies:
$\tilde{\Pi}^{*}_{ij}=(\alpha_{i}u_{i}^{1/c})\tilde{K}_{ij}(\beta_{j}v_{j}^{1/c}),$
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha u^{1/c})\tilde{K}\operatorname{diag}(\beta v^{1/c}),$ (g1)
$\implies\tilde{\Pi}^{*}=\operatorname{diag}(\alpha)\Pi^{1/c}\operatorname{diag}(\beta).$
Subject to the constraints:
$\sum_{i}\alpha_{i}\beta_{j}\Pi_{ij}^{*1/c}=a_{i},\quad\sum_{i}\alpha_{i}\beta_{j}\Pi_{ij}^{*1/c}=b_{j}.$
For any pair (i,j)&(k,l)(i,j)\&(k,l), we can express:
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}=\frac{\alpha_{i}}{\alpha_{k}}\frac{\beta_{j}}{\beta_{l}}\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
Taking logarithms on both sides, we have:
$\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)=\log(\alpha_{i})-\log(\alpha_{k})+\log(\beta_{j})-\log(\beta_{l})+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right).$
Let $\log(\alpha)=\phi$ and $\log(\beta)=\psi$, then:
$\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)=(\phi_{i}-\phi_{k})+(\psi_{j}-\psi_{l})+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right).$
This implies:
$\left|\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)-\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\right|\leq|\phi_{i}|+|\phi_{k}|+|\psi_{j}|+|\psi_{l}|.$
From Proposition 3 Proposition 2, we have:
$\max_{i}\phi_{i}\leq E,\quad\max_{i}\psi_{i}\leq E.$
Thus:
$\left|\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)-\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\right|\leq 4E.$
Therefore:
$4E+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)\leq\log\left(\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\right)\leq 4E+\frac{1}{c}\log\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right).$
This implies:
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\exp(4E)\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
Let $\gamma=\exp(4E)$, then:
$\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}\leq\gamma\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{1/c}.$
Corollary 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then:
$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s,$
where $m$ and $s$ are constants that depend on $\Pi^{*}$, the marginalization constants $a,b$ and the normalization constant $c$.
Proof.
From equation (g1) in Proposition 1 above, we know that:
$\tilde{\Pi}^{*}_{ij}=(\Pi^{*}_{ij})^{1/c}\cdot\exp(\phi_{i},\psi_{j})$
and from Proposition 2, we have that,
$\tilde{\Pi}^{*}_{ij}\leq(\Pi^{*}_{ij})^{1/c}\cdot e^{2E}$
$\Rightarrow\log(\tilde{\Pi}^{*}_{ij})\leq\frac{1}{c}\log(\Pi^{*}_{ij})+2E$
$\Rightarrow-\tilde{\Pi}^{*}_{ij}\log(\tilde{\Pi}^{*}_{ij})\geq-\frac{1}{c}(\Pi^{*}_{ij})^{1/c-1}\cdot\Pi^{*}_{ij}\log(\Pi^{*}_{ij})\cdot e^{2E}-2E\cdot e^{2E}\cdot(\Pi^{*}_{ij})^{1/c}$
For $c\gg 1$, $\frac{1}{c}\to 0$:
$\Rightarrow-\tilde{\Pi}^{*}_{ij}\log(\tilde{\Pi}^{*}_{ij})$ $\geq-\frac{1}{c\Pi^{*}_{ij}}\cdot\Pi^{*}_{ij}\log(\Pi^{*}_{ij})\cdot e^{2E}-2E\cdot e^{2E}\cdot(\Pi^{*}_{ij})^{1/c}$
$\Rightarrow-\tilde{\Pi}^{*}_{ij}\log(\tilde{\Pi}^{*}_{ij})$ $\geq-\frac{1}{c\Pi^{*}_{\min}}\cdot\Pi^{*}_{ij}\log(\Pi^{*}_{ij})\cdot e^{2E}-2E\cdot e^{2E}\cdot(\Pi^{*}_{ij})^{1/c}$
Summing for all $(i,j)$ we get,
$H(\tilde{\Pi^{*}})\geq mH(\Pi^{*})-s,$
where $m=\frac{e^{2E}}{c\Pi^{*}_{\min}}$ and $s=2E\cdot e^{2E}$. ∎
Proposition 2.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\Pi\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Consider the scaling factors $\alpha,\beta$ such that: $\tilde{u}_{i}=\alpha_{i}u_{i}^{1/c}$, $\tilde{v}_{j}=\beta_{j}v_{j}^{1/c}$ where $u,v$ are the Sinkhorn algorithm converged vectors for the original setting and $\tilde{u},\tilde{v}$ are for the cost-scaled version. Then, we have
$\max\{\left\|\phi\right\|_{\infty},\left\|\psi\right\|_{\infty}\}\leq\|M^{-1}\|_{\infty}\cdot\left\|\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}\right\|_{\infty}$,
where $\phi=\log(\alpha)$ and $\psi=\log(\beta)$. We also have that,
$\max_{i}|\alpha_{i}-1|,\max_{i}|\beta_{i}-1|\leq\|M^{-1}\|_{\infty}\max(\|\Delta_{a}\|_{\infty},\|\Delta_{b}\|_{\infty}),$
where $M,\Delta_{a},\Delta_{b}$ depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
Proof.
Let $X_{ij}=\Pi_{ij}^{*1/c}$ and $X=\Pi^{*1/c}$. Consider the exponentiated versions of $\alpha$ and $\beta$:
$\phi=\log(\alpha)\in\mathbb{R}^{n},\quad\psi=\log(\beta)\in\mathbb{R}^{m}.$
From the marginal constraints, we have:
$\sum_{j}X_{ij}e^{\phi_{i}+\psi_{j}}=a_{i},\quad\sum_{i}X_{ij}e^{\phi_{i}+\psi_{j}}=b_{j}.$
Applying a first-order Taylor expansion gives:
$\sum_{j}X_{ij}(1+\phi_{i}+\psi_{j})=a_{i}\quad\implies\quad\sum_{j}X_{ij}(\phi_{i}+\psi_{j})=a_{i}-\sum_{j}X_{ij},$
$\sum_{i}X_{ij}(1+\phi_{i}+\psi_{j})=b_{j}\quad\implies\quad\sum_{i}X_{ij}(\phi_{i}+\psi_{j})=b_{j}-\sum_{i}X_{ij}.$
Define:
$\Delta_{a_{i}}=a_{i}-\sum_{j}X_{ij},\quad\Delta_{b_{j}}=b_{j}-\sum_{i}X_{ij}.$
Thus, we have:
$\sum_{j}X_{ij}(\phi_{i}+\psi_{j})=\Delta a_{i},\quad\sum_{i}X_{ij}(\phi_{i}+\psi_{j})=\Delta b_{j}.$
This implies:
$\phi_{i}\left(\sum_{j}X_{ij}\right)+\sum_{j}X_{ij}\psi_{j}=\Delta a_{i},$
$\sum_{i}X_{ij}\phi_{i}+\psi_{j}\left(\sum_{i}X_{ij}\right)=\Delta b_{j}.$
Let:
$D_{r}=\operatorname{diag}(X\mathbf{1})\in\mathbb{R}^{n\times n},\quad D_{c}=\operatorname{diag}(X^{T}\mathbf{1})\in\mathbb{R}^{m\times m}.$
Then we can express the system as:
$\begin{pmatrix}D_{r}&X\\ X^{T}&D_{c}\end{pmatrix}\begin{pmatrix}\phi\\ \psi\end{pmatrix}=\begin{pmatrix}\Delta a\\ \Delta b\end{pmatrix}$.
Let:
$M=\begin{pmatrix}D_{r}&X\\ X^{T}&D_{c}\end{pmatrix}$.
Thus:
$\begin{pmatrix}\phi\\ \psi\end{pmatrix}=M^{-1}\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}$.
This implies:
$\left\|\begin{pmatrix}\phi\\ \psi\end{pmatrix}\right\|\leq\|M^{-1}\|\cdot\left\|\begin{pmatrix}\Delta_{a}\\ \Delta_{b}\end{pmatrix}\right\|$.
Since $\alpha=\exp(\phi)$ and $\beta=\exp(\psi)$, by assumption:
$|\alpha_{i}-1|\approx|\exp(\phi_{i})-1|\approx\phi_{i},$
$|\beta_{j}-1|\approx|\exp(\psi_{j})-1|\approx\psi_{j}.$
Therefore:
$\max_{i}|\alpha_{i}-1|,\max_{j}|\beta_{j}-1|\leq\|M^{-1}\|_{\infty}\cdot\max(\|\Delta a\|_{\infty},\|\Delta b\|_{\infty}).$
Theorem 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a general cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a prior transition probability matrix. Suppose $\Pi^{*}_{\mathrm{CTF-H}}$ is the solution to the following prior-aware optimal transport problem:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
where $\epsilon>0$ is the regularization parameter. Then, we can show that $\Pi^{*}_{\mathrm{CTF-H}}$ can be computed by the Sinkhorn algorithm and takes the form $\mathrm{diag}(\bm{u})\cdot\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)\cdot\mathrm{diag}(\bm{v})$, where $\odot$ stands for the elementwise multiplication, and $\bm{u}\in\mathbb{R}^{n_{0}},\bm{v}\in\mathbb{R}^{n_{1}}$ are vectors satisfying the marginalization constraints.
Proof.
We have that:
$\Pi^{*}_{\mathrm{CTF-H}}=\operatorname*{arg\,min}_{\Pi\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
Subject to:
$\Pi\mathbf{1}=a,\quad\Pi^{\top}\mathbf{1}=b.$
This formulation is a standard convex optimization setting with constraints. The Lagrangian of this setting is:
$\mathcal{L}(\Pi,f,g)=\sum_{k,l}C_{kl}\Pi_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\left(\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-1\right)-\sum_{k}f_{k}\left(\sum_{l}\Pi_{kl}-a_{k}\right)-\sum_{l}g_{l}\left(\sum_{k}\Pi_{kl}-b_{l}\right)$
Differentiating with respect to $\Pi_{kl},f_{k},g_{l}$, we get:
$\frac{\partial\mathcal{L}}{\partial\Pi_{kl}}=C_{kl}+\epsilon\log\left(\frac{\Pi_{kl}}{M_{kl}}\right)-f_{k}-g_{l}$
Setting the derivative to zero:
$\epsilon\log\left(\frac{\Pi^{*}_{kl}}{M_{kl}}\right)=f_{k}-C_{kl}-g_{l}$
$\implies\frac{\Pi^{*}_{kl}}{M_{kl}}=e^{\frac{f_{k}}{\epsilon}}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
$\implies\Pi^{*}_{kl}=e^{\frac{f_{k}}{\epsilon}}M_{kl}e^{-\frac{C_{kl}}{\epsilon}}e^{\frac{g_{l}}{\epsilon}}$
Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ such that:
$u_{k}=e^{\frac{f_{k}}{\epsilon}},\quad v_{l}=e^{\frac{g_{l}}{\epsilon}}$
Let $K_{kl}$ be the kernel $M_{kl}e^{-C_{kl}/\epsilon}$.
Then, we have:
$\Pi^{*}_{kl}=u_{k}K_{kl}v_{l}$
$\Pi^{*}=\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)$ (13)
Differentiating the Lagrangian with respect to $f_{k}$ and $g_{l}$, we get:
$\frac{\partial\mathcal{L}}{\partial f_{k}}=1\cdot\left(\sum_{l}\Pi^{*}_{kl}-a_{k}\right)=0$
$\implies\Pi^{*}\mathbf{1}=a$ (14)
$\frac{\partial\mathcal{L}}{\partial g_{l}}=1\cdot\left(\sum_{i}\Pi^{*}_{kl}-b_{l}\right)=0$
$\implies\Pi^{*\top}\mathbf{1}=b$ (15)
From equations 16 13, 17 14, and 18 15 above, we get:
$\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v)\cdot\mathbf{1}$ $=a$
$(\operatorname{diag}(u)\cdot K\cdot\operatorname{diag}(v))^{\top}\mathbf{1}$ $=b$
Which can be rewritten as:
$u\odot(Kv)$ $=a$
$K^{\top}u\odot v$ $=b$
This is the usual matrix scaling formulation for which the Iterative Proportional Fitting (IPF) updates are:
$u_{k}^{t+1}=\frac{a_{k}}{(Kv^{t})_{k}},\quad v_{l}^{t+1}=\frac{b_{l}}{(K^{\top}u^{t+1})_{l}}$
Sinkhorn Algorithm uses these updates, iteratively, and these updates are shown to converge in Franklin & Lorenz (1989). Thus, Sinkhorn Algorithm can be used for the ContextFlow’s Prior Aware Entropy Regularized (PAER) (CTF-H) formulation.
From equation (9) 13, we get:
$\Pi^{*}_{kl}=e^{f_{k}/\epsilon}M_{kl}e^{-C_{kl}/\epsilon}e^{g_{l}/\epsilon}$
When $\epsilon\to\infty$, we have $C_{kl}/\epsilon\to 0$.
$\quad e^{-C_{kl}/\epsilon}\to 1$
$\implies$ $\Pi^{*}_{kl}\to u_{k}M_{kl}v_{l}$
$\implies$ $\Pi^{*}_{\text{CTF-H}}\to\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$
Such that marginal constraints, $\Pi^{*}_{\text{CTF-H}}\mathbf{1}=a$ and $\Pi^{*\top}_{\text{CTF-H}}\mathbf{1}=b$ are satisfied. ∎
Appendix C Effects of normalization on Prior aware cost matrix
From Peyré et al. (2019), we know that optimal MOTFM coupling takes the form $\Pi^{*}_{\text{EOT}}=\text{diag}(u)\cdot K\cdot\text{diag}(v)$, where $K$ is the kernel matrix such that $K]_{ij}=\exp(\frac{-c_{ij}}{\epsilon})$, with $u,v$ satisfying marginalization constraints $u\odot Kv=a$ and $K^{T}u\odot v=b$. Sinkhorn updates are given by:
$u^{l+1}=\frac{a}{Kv^{l}};v^{l+1}=\frac{b}{K^{T}u^{l+1}}.$
In cases where the OT cost function consists of information from different modalities the distances are usually normalized to have distances of a similar scale. Normalizing the cost results $\tilde{c}_{ij}=\frac{c_{ij}}{\epsilon}$ such that the new kernel matrix $K_{\text{norm}}]_{ij}=\exp(\frac{-c_{ij}}{C_{\text{max}}\epsilon})$ can cause numerical issues if $C_{\text{max}}\gg 1$. The cost normalization should be performed mindfully, when considering different pairwise distances, as in PACM Section 3. Intuitively, scaling the cost has the same effect as that of increasing $\epsilon$, making solutions more diffused.
Proposition 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\Pi^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right).$
Let $\tilde{\Pi}^{*}$ be the EOT-coupling where the cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then, for any indices $(i,j)$ and $(k,l)$ we have
$\frac{\widetilde{\Pi}^{*}_{ij}}{\widetilde{\Pi}^{*}_{kl}}\leq\gamma\,\left(\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}\right)^{\frac{1}{c}},$
where $\gamma$ depends on $\Pi^{*}_{ij},c$ and OT marginal constraints $a,b$.
From Proposition 1, let $\frac{\Pi^{*}_{ij}}{\Pi^{*}_{kl}}=m$, such that $m>1$ ($\Pi^{*}_{ij}>\Pi^{*}_{kl}$ or entries are faraway) then, for $c>1$, we have $\frac{\tilde{\Pi}^{*}_{ij}}{\tilde{\Pi}^{*}_{kl}}<m^{\frac{1}{c}}<m$, for $\gamma<1$, implying that faraway entries are squeezed together. This results in bringing probabilities that are far apart closer to each other or, in essence, in creating more diffused and less sharp couplings.
Corollary 1.
Let $\mathbf{C}\in\mathbb{R}^{n_{0}\times n_{1}}$ be a cost matrix and $\mathbf{M}\in\mathbb{R}^{n_{0}\times n_{1}}$ a prior transition matrix with positive entries. Consider the entropy-regularized OT formulation:
$\mathbf{\Pi}^{*}$ $={\arg\!\min}_{\mathbf{\Pi}\geq 0}\;\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log\!\left(\Pi_{kl}\right)$
and $\tilde{\Pi^{*}}$ be EOT-coupling in the case when cost is scaled by a normalization constant $c$ or $\tilde{C}_{ij}=\frac{C_{ij}}{c}$. Let the regularization parameter $\epsilon>0$ be the same in both cases. Then we have:
$H(\tilde{\Pi}_{ij})\geq mH(\Pi_{ij})-s$
where $m$ and $s$ are constants, that depend on $\Pi^{*}$, marginalization constants $a,b$ and normalization constant $c$.
Corollary 1 can also be interpreted as supporting the results of Proposition 1 and our intuition that normalizing has the same effect on the kernel matrix as increasing $\epsilon$, leading to more diffused couplings or couplings with increased entropy.
Appendix D ContextFlow Algorithm
Algorithm 1 ContextFlow (CTF): Flow Matching with Spatial-Context-Aware OT Couplings
1:Input: gene data $\{\mathbf{X}_{t_{1}},\cdots,\mathbf{X}_{t_{m+1}}\}$, spatial data $\{\mathbf{S}_{t_{1}},\ldots,\mathbf{S}_{t_{m+1}}\}$, parameters $\lambda$, $\alpha$, $\epsilon$, $\sigma$, $\eta$
2:Output: neural velocity vector field $u_{\theta}$
3:Initialize $\theta$
4:while training do
5:  for $i=1,2,\ldots,m$ do
6:   Sample a batch $\mathcal{B}=\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim(\mathbf{X}_{t_{i}},\mathbf{X}_{t_{i+1}})\}$
7:   Construct TPM: $\mathbf{M}_{i,i+1}(\mathcal{B})\triangleright$ $\mathbf{M}_{i,i+1}$ is defined in Equation 10
8:   if “prior-aware cost matrix” then
9:     $C_{kl}\leftarrow\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot[\mathbf{M}_{i,i+1}]_{kl}$ for any pair $(k,l)$
10:     $\mathbf{K}\leftarrow\exp(-\mathbf{C}/\epsilon)$
11:   else if “prior-aware entropy regularization” then
12:     $C_{kl}\leftarrow\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ for any pair $(k,l)$
13:     $\mathbf{K}\leftarrow\widehat{\mathbf{M}}_{i,i+1}\odot\exp(-\mathbf{C}/\epsilon)$ $\triangleright$ $\widehat{\mathbf{M}}_{i,i+1}$ is defined in Equation 12
14:   end if
15:   Initialize $\bm{a}\leftarrow\frac{1}{n_{i}}\mathbf{1}_{n_{i}}$, $\bm{b}\leftarrow\frac{1}{n_{i+1}}\mathbf{1}_{n_{i+1}}$, $\bm{u}\leftarrow\mathbf{1}_{n_{i}}$, $\bm{v}\leftarrow\mathbf{1}_{n_{i+1}}$
16:   while not converged do
17:     $\bm{u}\leftarrow\bm{a}\oslash(\mathbf{K}\bm{v}),\>\bm{v}\leftarrow\bm{b}\oslash(\mathbf{K}^{\top}\bm{u})$ $\triangleright$ Run Sinkhorn algorithm
18:   end while
19:   Obtain spatial-prior-aware OT couplings $\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\leftarrow\operatorname{diag}(\bm{u})\mathbf{K}\operatorname{diag}(\bm{v})$
20:   Sample $t\sim\mathcal{U}(t_{i},t_{i+1})$ and $\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\}$
21:   Sample $\bm{x}_{t}\sim\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right)$
22:   $L_{\mathrm{CFM}}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{t,(\bm{x}_{i},\bm{x}_{i+1})}\left\|u_{\theta}(\bm{x}_{t},t)-\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\right\|_{2}^{2}$
23:  end for
24:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}L_{\mathrm{CFM}}$
25:end while
Appendix E Kantarovich-OT Formulation
Kantorovich’s formulation (Peyré et al., 2019) is a classical definition of the optimal transport (OT) problem that seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the 2-Wasserstein distance:
$\pi^{*}_{\mathrm{ot}}:={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (16)
where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$. Equation 16 can be solved in a mini-batch fashion using standard solvers such as POT (Flamary et al., 2021); however, the computational complexity is cubic in batch size.
Appendix F Evaluation Metrics
F.1 2-Wasserstein
The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
$\mathcal{W}_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\bigg(\sum_{(\bm{x},\bm{y})}\gamma(\bm{x},\bm{y})\cdot\|\bm{x}-\bm{y}\|_{2}^{2}\bigg)^{1/2},$
where $\Pi(\mu,\nu)$ denotes the set of couplings between $\mu$ and $\nu$.
F.2 Weighted 2-Wasserstein
Implausible velocity fields can steer a cell’s transcriptional trajectory in unrealistic directions, potentially leading to entirely different terminal cell types. We thus employ the weighted 2-Wasserstein metric, which ensures the evaluation accounts for both transcriptional similarity and the distributional balance of cell types. We define the weighted 2-Wasserstein distance (Weighted $\mathcal{W}_{2}$) between true and predicted distributions as:
$\text{Weighted-$\mathcal{W}_{2}$}(\mu,\nu)=\sum_{i=1}^{C}\frac{n_{i}^{\text{true}}}{N}\cdot\mathcal{W}_{2}\!\bigg(\frac{1}{n_{i}^{\text{true}}}\!\!\sum_{j:y_{j}=i}\delta_{\bm{x}_{j}},\frac{1}{n_{i}^{\text{pred}}}\!\!\sum_{j:\hat{y}_{j}=i}\delta_{\bm{x}_{j}}\bigg),$
where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples. To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen & Guestrin, 2016) trained for each dataset.
F.3 Energy Distance
Let $\mu$ and $\nu$ be probability distributions with samples $X=\{\bm{x}_{i}\}_{i=1}^{m}\sim\mu$ and $Y=\{\bm{y}_{j}\}_{j=1}^{n}\sim\nu$. The squared empirical energy distance (Energy) is defined as:
$\mathrm{ED}(\mu,\nu)=\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\|\bm{x}_{i}-\bm{y}_{j}\|-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{i^{\prime}=1}^{m}\|\bm{x}_{i}-\bm{x}_{i^{\prime}}\|-\frac{1}{n^{2}}\sum_{j=1}^{n}\sum_{j^{\prime}=1}^{n}\|\bm{y}_{j}-\bm{y}_{j^{\prime}}\|,$
where $\|\cdot\|$ is the Euclidean norm. The distance is non-negative and equals zero if and only if $\mu=\nu$.
F.4 Maximum Mean Discrepancy
For the same samples, the unbiased empirical estimate of the squared maximum mean discrepancy (MMD) with kernel $\kappa$ is defined as:
$\mathrm{MMD}(\mu,\nu;\kappa)=\frac{1}{m(m-1)}\sum_{i\neq i^{\prime}}\kappa(\bm{x}_{i},\bm{x}_{i^{\prime}})+\frac{1}{n(n-1)}\sum_{j\neq j^{\prime}}\kappa(\bm{y}_{j},\bm{y}_{j^{\prime}})-\frac{2}{mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\kappa(\bm{x}_{i},\bm{y}_{j}).$
In our evaluations, we use a multi-kernel variant with radial basis function (RBF) kernels $\kappa_{\gamma}(\bm{x},y)=\exp(-\gamma\|\bm{x}-\bm{y}\|^{2})$, and average over $\gamma\in[2,1,0.5,0.1,0.01,0.005]$.
Appendix G Biological Priors and Dataset Visualizations
G.1 (Im-)Plausibility of OT-Couplings
To demonstrate the need for integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework. From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types. Figures 3(a) and 3(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings. Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
In our transport plan couplings, we observed the following cell type lineage switches:
  • Immature MSN $\rightarrow$ Immature nptxEX
  • Immature MSN $\rightarrow$ Immature dpEX
  • Immature MSN $\rightarrow$ Immature CMPN
  • Immature nptxEX $\rightarrow$ Immature cckIN
  • Immature nptxEX $\rightarrow$ Immature MSN
(a) MOTFM Implausible Transitions
(b) ContextFlow Implausible Transitions
Figure 3: Comparison of biologically implausible cell type couplings between Stage 0 and Stage 1 of the Brain Regeneration Dataset (Wei et al., 2022), under the Entropic-OT and ContextFlow Regularized-OT formulations. Biological implausibility is defined here as transitions involving excitatory–inhibitory lineage switches. Our formulation produces substantially fewer biologically implausible couplings (24) compared to MOTFM (54).
Of these, $54$ implausible transitions arose from the Entropic-OT plan compared to the $24$ under the PAER-OT plan, with the specific transitions detailed in the figure legends. We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right. In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information. These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
G.2 Cell type distributions over time
Figures 46 present the spatial maps of the transcriptomics datasets across different time points, illustrating how tissue organization and cell type distributions evolve during development and regeneration. These maps highlight not only changes in cellular composition but also the preservation of spatial neighborhoods and geometrical arrangements of specific cell types over time. Such contextual information, specific to spatial transcriptomics, remains inaccessible to standard flow-matching frameworks. By contrast, ContextFlow is designed to exploit these spatial features, enabling the inference of trajectories that are both temporally smooth and spatially coherent.
G.2.1 Brain Regeneration
Figure 4: Temporal progression of spatial distribution of different cell types for Brain Regeneration.
G.2.2 Mouse Embryo Organogenesis
Figure 5: Temporal progression of spatial distribution of different cell types for Mouse Organogenesis.
G.2.3 Liver Regeneration
Figure 6: Temporal progression of spatial distribution of fibrogenic states for Liver Regeneration. Here, $0/1$ refers to the absence or presence of fibrogenic spots.
G.3 Ligand Receptor Interactions
Figure 7 shows the ligand-receptor score of the NPTX2-NPTXR pair in two consecutive slides from the Brain regeneration dataset (Wei et al., 2022). Similar activities are visible bilaterally in the cerebral cortex, suggesting that ligand–receptor interactions are preserved across time and spatially aligned with underlying tissue structure. This observation provides strong evidence that including LR interactions as contextual priors is biologically meaningful, as they capture functional communication signals between cells that remain stable across short time intervals.
(a) NPTX2-NPTXR LR pair activation on Stage 3
(b) NPTX2-NPTXR LR pair activation on Stage 4
Figure 7: Spatial distributions of LR activation for NPTX2-NPTXR in two consecutive slides from the Brain regeneration dataset. Similar activations are visible at structurally equal positions.
Based on the activation of NPTX2–NPTXR in Figure 7, we observe that the corresponding communication pattern naturally biases the optimal couplings towards transitions such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 8). These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context. This example highlights the richness of the contextual information captured by our proposed biological prior, and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process leads to more interpretable and biologically consistent trajectories.
Figure 8: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for the two consecutive slides.
Appendix H Additional Experiments & Ablations
H.1 Next Step Sampling for Axolotl Brain Regeneration
Table 5: Interpolation via Next Step Sampling at holdout time 3 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $2.618\pm 0.142$ $2.579\pm 0.197$ $0.043\pm 0.003$ $12.505\pm 1.271$
MOTFM $2.567\pm 0.088$ $2.476\pm 0.161$ $0.040\pm 0.003$ $11.269\pm 1.388$
CTF-C 11 $0.2$ $2.503\pm 0.071$ $2.425\pm 0.239$ $0.037\pm 0.003$ $9.868\pm 1.293$
11 $0.5$ $2.467\pm 0.107$ $2.301\pm 0.163$ $0.037\pm 0.002$ $9.532\pm 1.093$
11 $0.8$ $423\pm 0.164 2.2$$293\pm 0.103$ $0.037\pm 0.001$ $9.874\pm 0.659$
0 $0.2$ $2.396\pm 0.028$ $2.100\pm 0.102$ $0.033\pm 0.003$ $8.577\pm 0.976$
0 $0.5$ $2.447\pm 0.142$ $2.337\pm 0.216$ $0.036\pm 0.005$ $9.696\pm 1.882$
0 $0.8$ $413\pm 0.099 2.2$$293\pm 0.161$ $0.036\pm 0.002$ $9.114\pm 1.092$
$0.5$ $0.2$ $2.460\pm 0.118$ $2.342\pm 0.144$ $0.036\pm 0.003$ $9.500\pm 1.067$
$0.5$ $0.5$ $2.504\pm 0.094$ $2.309\pm 0.139$ $0.036\pm 0.003$ $9.394\pm 1.431$
$0.5$ $0.8$ $2.442\pm 0.173$ $2.353\pm 0.241$ $0.035\pm 0.004$ $9.008\pm 2.094$
CTF-H 0 $2.528\pm 0.143$ $2.534\pm 0.180$ $0.040\pm 0.004$ $11.192\pm 1.304$
11 $\bm{2.316\pm 0.141}$ $\bm{1.969\pm 0.221}$ $\bm{0.030\pm 0.004}$ $\bm{6.359\pm 1.336}$
$0.5$ $2.519\pm 0.167$ $2.412\pm 0.158$ $0.039\pm 0.004$ $10.304\pm 1.808$
Table 6: Extrapolation via Next Step Sampling at holdout time 5 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step CFM $7.124\pm 0.443$ $7.133\pm 0.533$ $0.276\pm 0.011$ $76.947\pm 5.661$
MOTFM $7.487\pm 0.698$ $7.449\pm 0.931$ $0.266\pm 0.010$ $81.965\pm 9.812$
CTF-C 11 $0.2$ $7.257\pm 0.597$ $7.077\pm 0.473$ $0.257\pm 0.004$ $79.562\pm 7.787$
11 $0.5$ $68\pm 0.608 6.969$$69\pm 0.628$ $0.265\pm 0.009$ $77.025\pm 6.056$
11 $0.8$ $695\pm 0.443 7.7$$792\pm 0.463$ $0.266\pm 0.007$ $87.179\pm 6.690$
0 $0.2$ $8.170\pm 0.663$ $8.079\pm 0.723$ $0.269\pm 0.008$ $91.572\pm 8.802$
0 $0.5$ $7.244\pm 0.804$ $7.146\pm 0.775$ $0.265\pm 0.003$ $80.424\pm 10.376$
0 $0.8$ $7.382\pm 1.068$ $7.234\pm 0.852$ $0.267\pm 0.009$ $81.635\pm 14.135$
$0.5$ $0.2$ $94\pm 0.239 7.171$$71\pm 0.422$ $0.266\pm 0.001$ $78.924\pm 3.715$
$0.5$ $0.5$ $7.188\pm 0.391$ $\bm{6.931\pm 0.260}$ $0.267\pm 0.005$ $78.992\pm 6.195$
$0.5$ $0.8$ $7.242\pm 0.804$ $7.166\pm 0.980$ $0.267\pm 0.006$ $80.509\pm 10.304$
CTF-H 0 $\bm{6.914\pm 0.471}$ $7.198\pm 0.726$ $0.266\pm 0.009$ $\bm{76.149\pm 8.436}$
11 $7.505\pm 0.667$ $7.338\pm 0.601$ $\bm{0.263\pm 0.006}$ $83.425\pm 8.793$
$0.5$ $7.243\pm 0.479$ $7.157\pm 0.641$ $0.270\pm 0.007$ $79.826\pm 8.067$
H.2 IVP Sampling on Axolotl Brain Regeneration
Table 7: Interpolation via IVP Sampling at time point 3 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP CFM $4.216\pm 0.463$ $4.266\pm 0.308$ $0.170\pm 0.029$ $32.413\pm 5.122$
MOTFM $198\pm 0.319 4.4$$452\pm 0.243$ $0.173\pm 0.017$ $33.149\pm 3.321$
CTF-C 11 $0.2$ $4.011\pm 0.276$ $4.048\pm 0.321$ $0.147\pm 0.021$ $30.337\pm 4.713$
11 $0.5$ $3.932\pm 0.377$ $4.356\pm 0.398$ $0.156\pm 0.025$ $31.524\pm 4.875$
11 $0.8$ $3.603\pm 0.300$ $3.816\pm 0.310$ $0.127\pm 0.018$ $24.271\pm 3.992$
0 $0.2$ $\bm{3.465\pm 0.232}$ $\bm{3.641\pm 0.320}$ $0.119\pm 0.025$ $23.055\pm 5.939$
0 $0.5$ $3.943\pm 0.413$ $4.241\pm 0.435$ $0.150\pm 0.039$ $29.221\pm 5.713$
0 $0.8$ $3.881\pm 0.368$ $4.094\pm 0.551$ $0.139\pm 0.026$ $27.941\pm 6.676$
$0.5$ $0.2$ $4.152\pm 0.341$ $4.322\pm 0.291$ $0.166\pm 0.014$ $33.299\pm 3.629$
$0.5$ $0.5$ $4.013\pm 0.187$ $4.138\pm 0.297$ $0.153\pm 0.020$ $30.941\pm 3.685$
$0.5$ $0.8$ $4.015\pm 0.351$ $3.974\pm 0.442$ $0.140\pm 0.038$ $27.592\pm 6.669$
CTF-H 0 $3.925\pm 0.267$ $4.375\pm 0.297$ $0.164\pm 0.013$ $32.034\pm 3.270$
11 $3.905\pm 0.395$ $4.188\pm 0.685$ $\bm{0.074\pm 0.014}$ $\bm{18.728\pm 2.689}$
$0.5$ $3.917\pm 0.343$ $4.159\pm 0.455$ $0.147\pm 0.022$ $29.613\pm 4.822$
Table 8: Extrapolation via IVP Sampling at holdout time 5 for the Brain Regeneration dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP CFM $6.633\pm 1.312$ $7.116\pm 1.084$ $0.143\pm 0.037$ $60.573\pm 21.756$
MOTFM $6.503\pm 0.720$ $6.352\pm 0.592$ $0.162\pm 0.038$ $56.452\pm 15.932$
CTF-C 11 $0.2$ $6.403\pm 0.959$ $6.558\pm 1.297$ $0.160\pm 0.024$ $61.051\pm 16.594$
11 $0.5$ $6.260\pm 0.616$ $7.681\pm 4.003$ $0.157\pm 0.039$ $52.478\pm 12.010$
11 $0.8$ $6.875\pm 0.643$ $6.920\pm 0.796$ $0.159\pm 0.045$ $62.838\pm 16.897$
0 $0.2$ $6.722\pm 0.905$ $6.782\pm 1.003$ $0.154\pm 0.034$ $53.996\pm 15.617$
0 $0.5$ $6.614\pm 0.710$ $6.854\pm 0.740$ $0.201\pm 0.023$ $70.370\pm 9.099$
0 $0.8$ $6.504\pm 0.925$ $6.744\pm 1.336$ $0.174\pm 0.037$ $56.687\pm 18.118$
$0.5$ $0.2$ $6.514\pm 0.504$ $5.998\pm 0.803$ $0.155\pm 0.032$ $51.329\pm 15.080$
$0.5$ $0.5$ $6.696\pm 0.427$ $6.481\pm 0.387$ $0.195\pm 0.024$ $66.212\pm 3.542$
$0.5$ $0.8$ $6.550\pm 0.975$ $6.563\pm 1.029$ $0.188\pm 0.037$ $63.014\pm 14.173$
CTF-H 0 $6.243\pm 0.760$ $6.220\pm 0.751$ $0.195\pm 0.020$ $61.316\pm 10.288$
11 $\bm{5.277\pm 0.936}$ $6.021\pm 1.192$ $\bm{0.099\pm 0.007}$ $\bm{27.777\pm 8.621}$
$0.5$ $6.254\pm 0.819$ $\bm{5.973\pm 0.757}$ $0.156\pm 0.025$ $54.330\pm 12.089$
H.3 Next Step Sampling for Mouse Embryo Organogenesis
Table 9: Interpolation via Next Step Sampling at holdout time 5 for the Mouse Organogenesis dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step MOTFM $1.892\pm 0.028$ $1.873\pm 0.086$ $0.164\pm 0.002$ $11.615\pm 0.092$
CTF-C 11 $0.2$ $1.881\pm 0.020$ $1.922\pm 0.078$ $0.158\pm 0.003$ $11.529\pm 0.197$
11 $0.5$ $\bm{1.865\pm 0.030}$ $1.852\pm 0.093$ $0.159\pm 0.001$ $11.482\pm 0.108$
11 $0.8$ $1.889\pm 0.024$ $1.888\pm 0.082$ $0.161\pm 0.002$ $11.552\pm 0.166$
0 $0.2$ $1.893\pm 0.035$ $1.912\pm 0.057$ $0.159\pm 0.001$ $11.462\pm 0.154$
0 $0.5$ $1.877\pm 0.039$ $1.933\pm 0.088$ $0.162\pm 0.002$ $11.528\pm 0.110$
0 $0.8$ $1.882\pm 0.022$ $1.869\pm 0.049$ $0.161\pm 0.001$ $\bm{11.399\pm 0.119}$
$0.5$ $0.2$ $1.886\pm 0.022$ $1.927\pm 0.111$ $\bm{0.157\pm 0.002}$ $11.430\pm 0.131$
$0.5$ $0.5$ $1.899\pm 0.027$ $1.899\pm 0.072$ $0.160\pm 0.002$ $11.517\pm 0.097$
$0.5$ $0.8$ $1.888\pm 0.033$ $\bm{1.839\pm 0.134}$ $0.161\pm 0.002$ $11.475\pm 0.159$
CTF-H 0 $1.884\pm 0.027$ $1.862\pm 0.123$ $0.164\pm 0.001$ $11.499\pm 0.123$
11 $1.898\pm 0.029$ $1.866\pm 0.097$ $0.167\pm 0.002$ $11.795\pm 0.170$
$0.5$ $1.871\pm 0.030$ $1.919\pm 0.067$ $0.164\pm 0.002$ $11.639\pm 0.182$
Table 10: Extrapolation via Next Step Sampling at holdout time 8 for Mouse Organogenesis.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
Next Step MOTFM $1.626\pm 0.066$ $1.682\pm 0.096$ $0.084\pm 0.007$ $7.418\pm 0.749$
CTF-C 11 $0.2$ $1.683\pm 0.058$ $1.803\pm 0.117$ $0.087\pm 0.006$ $7.830\pm 0.551$
11 $0.5$ $1.685\pm 0.096$ $1.714\pm 0.159$ $0.089\pm 0.006$ $8.056\pm 1.033$
11 $0.8$ $1.703\pm 0.063$ $1.830\pm 0.131$ $0.095\pm 0.005$ $8.928\pm 0.723$
0 $0.2$ $1.715\pm 0.123$ $1.860\pm 0.267$ $0.094\pm 0.009$ $9.021\pm 1.740$
0 $0.5$ $1.725\pm 0.082$ $1.856\pm 0.191$ $0.093\pm 0.006$ $8.806\pm 0.749$
0 $0.8$ $1.774\pm 0.053$ $1.897\pm 0.175$ $0.094\pm 0.007$ $9.466\pm 0.957$
$0.5$ $0.2$ $1.818\pm 0.096$ $2.089\pm 0.222$ $0.084\pm 0.008$ $8.875\pm 0.976$
$0.5$ $0.5$ $1.774\pm 0.104$ $1.899\pm 0.280$ $0.093\pm 0.007$ $9.139\pm 1.437$
$0.5$ $0.8$ $1.768\pm 0.058$ $1.858\pm 0.120$ $0.101\pm 0.006$ $9.303\pm 0.634$
CTF-H 0 $\bm{1.505\pm 0.057}$ $\bm{1.397\pm 0.088}$ $0.087\pm 0.005$ $\bm{5.954\pm 0.492}$
11 $1.890\pm 0.046$ $1.877\pm 0.103$ $0.147\pm 0.006$ $10.752\pm 0.405$
$0.5$ $1.636\pm 0.060$ $1.684\pm 0.099$ $\bm{0.081\pm 0.005}$ $7.088\pm 0.692$
H.4 IVP Sampling for Mouse Embryo Organogenesis
Extrapolating to the last holdout time point of the mouse organogenesis dataset (Chen et al., 2022), particularly under IVP-Sampling, represents the most challenging setting among all our experiments. This difficulty arises because the target time point lies entirely outside the training horizon, requiring integration from the initial samples through to the end. As a result, the velocity field has more opportunity to drift in incorrect directions, often leading to generations that deviate substantially from the true dynamics. In our experiments, this instability was evident: across 10 runs, several produced highly unstable trajectories, reflecting the sensitivity of the system to initial conditions and numerical solvers. This variability is also captured in the performance metrics reported in Table 12.
Table 11: Interpolation via IVP Sampling at holdout time 5 for the Mouse Organogenesis dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP MOTFM $3.251\pm 0.676$ $3.418\pm 0.727$ $0.090\pm 0.003$ $9.226\pm 0.648$
CTF-C 11 $0.2$ $3.261\pm 0.880$ $5.264\pm 3.060$ $0.089\pm 0.003$ $10.724\pm 1.288$
11 $0.5$ $3.137\pm 0.407$ $4.093\pm 1.187$ $0.086\pm 0.004$ $11.948\pm 1.393$
11 $0.8$ $3.392\pm 0.757$ $4.716\pm 2.079$ $0.089\pm 0.005$ $9.547\pm 0.752$
0 $0.2$ $2.953\pm 0.425$ $3.816\pm 0.973$ $0.083\pm 0.002$ $9.816\pm 0.715$
0 $0.5$ $2.938\pm 0.476$ $3.904\pm 1.120$ $0.088\pm 0.005$ $9.864\pm 0.764$
0 $0.8$ $3.101\pm 0.539$ $3.855\pm 0.946$ $0.087\pm 0.004$ $9.280\pm 0.551$
$0.5$ $0.2$ $3.771\pm 0.862$ $5.457\pm 1.704$ $\bm{0.079\pm 0.004}$ $9.262\pm 1.134$
$0.5$ $0.5$ $3.090\pm 0.635$ $4.596\pm 2.357$ $0.084\pm 0.005$ $9.786\pm 1.067$
$0.5$ $0.8$ $3.200\pm 0.403$ $3.555\pm 0.637$ $0.084\pm 0.004$ $9.269\pm 0.541$
CTF-H 0 $3.244\pm 0.713$ $3.946\pm 1.671$ $0.089\pm 0.005$ $\bm{8.797\pm 0.612}$
11 $\bm{2.814\pm 0.414}$ $\bm{3.233\pm 0.567}$ $0.093\pm 0.005$ $10.319\pm 0.817$
$0.5$ $5.200\pm 0.799$ $6.306\pm 1.037$ $0.123\pm 0.008$ $45.862\pm 13.765$
Table 12: Extrapolation via IVP Sampling at holdout time 8 for the Mouse Organogenesis dataset.
Sampling Method $\bm{\lambda}$ $\bm{\alpha}$ Weighted $\bm{\mathcal{W}_{2}}$ $\bm{\mathcal{W}_{2}}$ MMD Energy
IVP MOTFM $110835\pm 211671$ $1021005\pm 2063905$ $0.086\pm 0.002$ $14178\pm 29475$
CTF-C 11 $0.2$ $785586\pm 1318212$ $7598321\pm 13497483$ $0.088\pm 0.002$ $98199\pm 150412$
11 $0.5$ $2691\pm 3931$ $28480\pm 36483$ $0.087\pm 0.002$ $1632\pm 2090$
11 $0.8$ $2473\pm 3349$ $19537\pm 26306$ $0.087\pm 0.003$ $517\pm 616$
0 $0.2$ $1493\pm 2497$ $14563\pm 24858$ $0.087\pm 0.001$ $800\pm 1158$
0 $0.5$ $218018\pm 471298$ $1820788\pm 3994886$ $0.086\pm 0.001$ $2170\pm 4697$
0 $0.8$ $2736\pm 34766 11$$18089\pm 310135$ $0.084\pm 0.002$ $27013\pm 60065$
$0.5$ $0.2$ $8114720\pm 16270274$ $69458305\pm 140579849$ $0.088\pm 0.002$ $901074\pm 1775139$
$0.5$ $0.5$ $2414338\pm 6009993$ $23103811\pm 56863018$ $0.086\pm 0.001$ $261335\pm 663279$
$0.5$ $0.8$ $158\pm 3023 11$$1138\pm 30025$ $0.084\pm 0.002$ $445\pm 1085$
CTF-H 0 $353428\pm 952168$ $3011396\pm 8057131$ $0.095\pm 0.004$ $22990\pm 58936$
11 $\bm{15\pm 10}$ $\bm{53\pm 53}$ $0.098\pm 0.006$ $\bm{48\pm 32}$
$0.5$ $107889\pm 275882$ $994606\pm 2772756$ $0.087\pm 0.002$ $8875\pm 24264$
H.5 Liver Regeneration
Table 13: Wasserstein distances for different model configurations
Variant $\bm{\lambda}$ $\bm{\alpha}$ $\bm{\mathcal{W}_{2}}$
EOT $34.30348\pm 1.44797$
CTF-C 11 $0.2$ $34.44455\pm 1.19306$
CTF-C 11 $0.5$ $33.95671\pm 1.64415$
CTF-C 11 $0.8$ $34.62812\pm 0.98181$
CTF-C 0 $0.2$ $34.24147\pm 1.16930$
CTF-C 0 $0.5$ $32.74147\pm 1.86351$
CTF-C 0 $0.8$ $33.71729\pm 1.23057$
CTF-C $0.5$ $0.2$ $33.56646\pm 1.04376$
CTF-C $0.5$ $0.5$ $33.84199\pm 1.71408$
CTF-C $0.5$ $0.8$ $33.04534\pm 1.64399$
CTF-H 0 $\bm{32.68215\pm 1.47185}$
CTF-H 11 $33.48050\pm 1.00149$
CTF-H $0.5$ $33.41444\pm 0.99501$
H.6 IVP Cell Type Progression over time
Figure 9: Temporal cell type predictions from ContextFlow for the major cell types in the Organogenesis dataset (Chen et al., 2022). Early progenitor populations (neural crest and mesenchyme) progressively diminish as development advances, while terminal fates (muscle, cartilage primordium, and liver) emerge at later stages. Major lineages such as brain, heart, and connective tissue remain continuous throughout. Overall, ContextFlow captures biologically coherent and temporally consistent developmental dynamics.